Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. Provide the solution to the system.
2x - 5y = -12
3x - 4y = 3

Answer by stanbon(75887)

use the elimination method to solve the following system of equations. Provide the solution to the system.
2x - 5y = -12
3x - 4y = 3
=======
To eliminate the "x" terms multiply the top equation by 3 and the
bottom equation by 2:
------
6x - 15y = -36
6x - 8y = 6
--------
Subtract the top from the bottom and solve for "y";
7y = 30
Divide both sides by 7:
y = 30/7
--------------------
Going back to the original equations to solve for "x":
2x - 5y = -12
3x - 4y = 3
Multiply thru the top equation by 4 and the bottom by 5:
8x - 20y = -48
15x - 20y = 15
---------------
Subtract the top equation from the bottom:
7x = 63
x = 9
-----------
Cheers,
Stan H.